# Mobile Edge Intelligence for Large Language Models: A Contemporary Survey

### v1

@@ II-A3 Multimodal LLMs
II-A4 Generative/Interactive AI
@@
II-A5 Industrial progress for LLMs
@@
II-B Mobile Edge Intelligence
MEI has emerged as a promising paradigm integrating AI with mobile edge computing, revolutionizing the landscape of mobile services and applications [70, 71, 72, 73]. The evolution of MEI comes from the convergence of various technological advancements, including the proliferation of Internet of Things (IoT) devices, the deployment of mobile networks, and the maturation of AI algorithms [70, 74, 75, 76]. These developments have enabled MEI to surmount the limitations of traditional cloud-centric architectures by providing localized AI training/inference and data processing capabilities at the network edge.
@@ II-B Mobile Edge Intelligence
On the application side, MEI carries substantial implications across various domains, such as smart healthcare, autonomous driving, and smart cities [80]. For instance, in the healthcare sector, MEI empowers real-time monitoring of patient health data and facilitates prompt interventions during emergencies. Likewise, in smart cities, MEI contributes to intelligent traffic management, environmental monitoring, and energy optimization, thereby fostering sustainability and enhancing the quality of life. Edge Intelligence has also witnessed notable industrial advancements, particularly with the proliferation of edge computing technologies and the advent of 5G networks. Leading enterprises, such as Microsoft, Google, Amazon, and NVIDIA, have developed edge AI platforms to support real-time AI services. For edge AI-empowered IoT applications, Microsoft “Azure IoT Edge”, Google “Cloud IoT”, Amazon “Web Services IoT”, and NVIDIA “EGX” provide edge AI platforms to bring real-time AI services across a wide range of applications, spanning from live video analytics [81], smart home [82], and industrial IoT [83].
II-C Lessons Learned for MEI4LLM
Clearly, MEI4LLM is nothing but a special case of MEI. However, the need to train and deploy a massive number of LLMs at the edge can be a key stimulus for the evolution of MEI. On the one hand, the principles in next-generation MEI, including the push for full integration of AI and communications, align well with what edge LLM needs. On the other hand, the extreme resource demands of LLMs push forward the boundary of MEI. Specifically, MEI4LLM must feature the following characteristics: 1) the native support of model splitting and parallel training/inference across interconnected edge nodes to facilitate the deployment of extremely large-scale models; 2) the integrated design of wireless networking and resource-efficient LLM training/inference techniques such as parameter-efficient fine-tuning and token (representation) reduction (which will be introduced in Section VII and Section VIII), to make LLM deployment cost-effective. In essence, compared to traditional MEI, MEI4LLM primarily focuses on exploring the integrated design of resource management and efficient AI techniques to support LLMs under limited communication-computing resources, which will be the focus of this survey paper and the research theme in this field.
III Preliminaries II: Resource-efficient LLM Techniques
Undoubtedly, due to the large sizes and computational complexity of LLMs, deploying LLMs on edge devices/servers for training/inference presents several critical challenges:
@@ III Preliminaries II: Resource-efficient LLM Techniques
  • Excessive computing overhead: It is reported that GPT-4 requires approximately 560 tera floating point operations per forward pass to generate one token [84]. However, the advanced A100 GPU provides a computational capacity of only 19.5 tera floating point operations per second [85]. This indicates that each forward pass of GPT-4 would take about 28 seconds to generate one token with a single A100 GPU. Furthermore, the backward propagation typically requires more computational resources than a forward pass [86], implying that on-device training would be even more challenging.
@@
  • Huge storage/memory demands: On the one hand, caching LLMs on edge devices consumes significant storage resources. LLMs specifically designed for on-device deployment even have billions of parameters, e.g., Google’s on-device Gemini Nano-2 has 3.25 billion parameters. On the other hand, optimizing the model with the commonly used Adam optimizer during training usually requires 12 times the memory resources needed for inference [87], which is unacceptable for mobile devices with limited memory. These factors indicate that deploying LLMs on edge devices for both training and inference poses stringent requirements on the storage and memory resources of edge devices.
@@
  • High energy costs: The limited volume of battery in edge devices hinders the deployment of LLMs on edge devices. For instance, running an LLM quantized to INT4 with 13 billion parameters using llama.cpp (one of the most lightweight on-device LLM engines) on a Xiaomi 11 smartphone results in an energy consumption of about 56 J/token [88]. This implies that a smartphone with a battery capacity of 3000 mAh and an output voltage of 3.7 V can generate only approximately 700 tokens if the LLM is deployed on the smartphone. The amount of data that can be processed would be even less if LLM training/fine-tuning were to be performed on edge devices.
@@
To mitigate the above challenges, in this section, we will review the enabling techniques tailored for resource-efficient LLM deployment, which are summarized in Fig. 3. The comparison of related works is shown in Table II. It is worth noting that the methods discussed in this section can decrease the complexity of LLM deployment on edge devices, edge servers, or in device-server collaboration. Consequently, these key techniques serve as the foundation for MEI4LLM and all subsequent sections.
@@
TABLE II: Summary of related works on resource-efficient LLM techniques.
@@
III-A Resource-efficient Inference
On-device LLM inference eliminates privacy leakage and the need for Internet connections. However, it poses significant challenges for on-device deployment since LLMs require substantial computing, memory, and energy resources. In what follows, we briefly present how to mitigate these problems by introducing efficient on-device LLM inference techniques.
III-A1 LLM compression
LLM compression enables the deployment of compressed LLMs on edge devices, significantly alleviating both memory/storage and computing demands on edge devices. The design objective is to compress LLMs without substantially compromising inference accuracy. Although compression has been extensively studied in the field of traditional DNNs, the unique architectures and properties of LLMs necessitate the redesign of existing compression strategies. The compression techniques tailored for LLMs are detailed below.
@@ III-A1 LLM compression
In LLM quantization, uneven weight quantization can be adopted to preserve the inference accuracy of quantized LLMs since not all weights in LLMs contribute equally to the final inference results [90]. For example, the authors in [90] propose the activation-aware weight quantization for uneven weight quantization. This method quantizes most of the weights in LLMs from FP16 to INT3/INT4. However, it retains approximately 1% of the critical weights in FP16 format, which significantly impacts quantization error.
@@
III-A2 Fast decoding for LLMs
Fast decoding methods can be used in on-device LLM deployment to save computing resources of edge devices in inference and facilitate on-device inference. Typical fast decoding methods can be categorized into speculative decoding, early exit, mixture-of-experts, contextual sparsity prediction, and parallel decoding, which are elaborated below.
@@ III-A2 Fast decoding for LLMs
Early exit: To minimize computing latency, the early exit strategy can be employed to bypass certain subsequent layers that input tokens traverse, effectively accelerating inference time. In the case of Transformers, early exit modules may be incorporated following some early blocks [96, 123], where the output from intermediate layers is converted into final outputs at exit points, provided that the desired confidence value is attained. However, the conventional early exit techniques may not be effective in LLM inference since the hidden states of the token, which output the results in an early exit layer, may be missing in the later layers [124]. For example, upon early exit, the hidden states and the key-value (KV) caches of the corresponding token in the subsequent layers will be missing, hindering the generation of future token sequences [125]. To address this challenge, the hidden states of the current token can be forwarded from the exiting layer to subsequent layers for KV cache computation [97, 126] if the token outputs the result in the exiting layer.
@@
III-B Resource-efficient Fine-tuning
@@
III-B1 Parameter-efficient fine-tuning
@@
III-B2 Zeroth-order optimization
Zeroth-order optimization [142] is a novel technique for model training, which estimates gradient updates only through the forward pass. This method considerably reduces the computational burden since the forward pass, equivalent to inference, demands far fewer computing resources than the backpropagation during training. Specifically, compared with prevalent first-order optimizers, such as Adam, zeroth-order optimizers do not need to store the intermediate results of the backpropagation during the training process, significantly reducing the memory consumption in LLM training. For example, the authors in [87] propose the zeroth-order optimizer, MeZO, which adopts simultaneous perturbation stochastic approximation to estimate model gradients only through the forward propagation and uses the estimated gradients to update model parameters. Compared with the fine-tuning with Adam, the model fine-tuned with MeZO demonstrates competitive performance on 7 out of 11 tasks, using only 1/12 of the running memory while only causing less than a 1% accuracy degradation. Moreover, to further enhance the efficiency of LLM fine-tuning, the zeroth-order optimization technique can be combined with PEFT methods, such as LoRA and prefix fine-tuning [87].
IV Application Scenarios
Although LLMs can be applied to a broad range of tasks, we focus on the application scenarios that motivate the deployment of LLMs at the network edge. Thus, as shown in Fig. 5, we present four killer LLM-empowered applications while concentrating on three aspects: latency requirements, bandwidth costs, and privacy requirements.
Latency reqs. $\leq$400 ms [143] 10-100 ms [79] 200 ms [144] 10 ms [79]
@@ V An Overview of MEI4LLM
Pushing LLMs to the network edge is a continuing trend with the killer applications in Section IV and the enabling resource-efficient techniques as introduced in III. By deploying LLMs at the network edge, edge devices can collaborate with edge servers for collaborative learning and inference, which largely mitigates the resource scarcity of edge devices. In line with the “NET4AI” (network for AI) vision for the 6G era [159], this section presents an overview of the MEI framework that supports the deployment of LLMs, called MEI4LLM, as shown in Fig. 6. MEI4LLM consists of the following basic components.
@@
Figure 6: The MEC architecture for LLMs in 6G with four basic components, i.e., AI-native architecture, parameter-sharing LLM caching and delivery, distributed LLM training (fine-tuning), and distributed LLM inference.
V-A AI-native Architecture
The next-generation edge networks will support AI services in an E2E manner. The goal of 6G should be to support AI, including LLMs, with superior performance with minimal communication, computing, storage, and energy requirements. For this reason, “task-oriented” architecture is commonly envisioned for 6G. Instead of maximizing throughput or minimizing latency, the design goal can be minimizing the cross entropy of output tokens from LLMs by implementing the optimal distributed computing, feature extraction, and resource allocation schemes under diverse resource constraints.
@@ V-A AI-native Architecture
To achieve this goal, network virtualization is of paramount importance for improving resource utilization, flexibility, and manageability. Following the design principle of software-defined networking, the MEI4LLM features a central controller that orchestrates network-wide computing resources and data transmissions, with the decoupled control and data plane. By collecting global network knowledge, such as the accuracy of LLMs, various quantization levels, user requirements on LLM services, channel conditions, battery status of users, and computing resource availability, the control partitions and coordinates model training/inference and delivery across the distributed edge computing systems, with intermediate smashed data (i.e., intermediate activations and back-propagated gradients), model parameters, or user data exchanged across edge routers and servers.
@@
Further edge networks will evolve into the “neural edge” [159], where neural network layers are distributed across edge nodes for collaborative computing. Analogous to cloud data centers with many GPUs to support large-scale LLMs, MET4LLM must feature flexible and model splitting for training and inference across distributed edge devices and servers. Both air interfaces and network designs should have native support for federated learning, split learning, and split inference for AI models, including LLMs. Since model training and inference are robust to packet errors, task-oriented wireless transmissions, say, for smashed data at the cut layer, can be carried out with appropriate error control to achieve the best efficiency-reliability trade-off. The optimal model splitting, placement, and data routing for large-scale models should be concertedly supported over edge networks.
@@
At last, information-centric networking can be implemented to ensure seamless model, feature, and data transmission across edge networks for efficient delivery of LLMs. In this respect, MEI4LLM should support LLM parameter block naming and name-based transmission protocol. By assigning names for each LLM parameter block, the central controller in the MEI4LLM architecture can forward the parameter request to where it is cached, reducing latency and bandwidth consumption for delivering large-scale models across networks and to end users.

### v2

@@ III-A3 Multimodal LLMs
III-A4 Practical applications: Generative/Interactive AI
@@
III-A5 Industrial progress for LLMs
@@
III-B Mobile Edge Intelligence
MEI has emerged as a promising paradigm integrating AI with mobile edge computing, revolutionizing the landscape of mobile services and applications [111, 112, 113, 114]. As the precursor to MEI, mobile edge computing allows connected edge servers in mobile edge networks, such as base stations and wireless access points, to provide computing services to resource-limited edge devices. This significantly reduces the computing burden on edge devices and avoids the need to directly transfer massive amounts of data to the cloud, which would otherwise incur long transmission latency [115, 116]. MEI can be viewed as the confluence of AI and mobile edge computing. By leveraging mobile edge computing technology, MEI enables connected edge servers to execute AI algorithms with end-edge-cloud synergy, facilitating data collection, caching, processing, and analysis for data captured by edge devices in mobile AI applications. This greatly enhances both the efficiency and speed of data processing while enhancing data privacy [117, 118, 119]. Specifically, MEI surmounts the limitations of traditional cloud-centric architectures by coordinating computing resources across a hierarchical structure that includes edge devices, edge servers, and cloud centers. This coordination can substantially optimize the performance of localized AI services, such as AI model caching, training, and inference [120, 118], thereby facilitating the deployment of intelligent applications at the network edge.
@@ III-B Mobile Edge Intelligence
On the application side, MEI carries substantial implications across various domains, such as smart healthcare, autonomous driving, and smart cities [123]. For instance, in the healthcare sector, MEI empowers real-time monitoring of patient health data and facilitates prompt interventions during emergencies. Likewise, in smart cities, MEI contributes to intelligent traffic management, environmental monitoring, and energy optimization, thereby fostering sustainability and enhancing the quality of life. Edge Intelligence has also witnessed notable industrial advancements, particularly with the proliferation of edge computing technologies and the advent of 5G networks. Leading enterprises, such as Microsoft, Google, Amazon, and NVIDIA, have developed edge AI platforms to support real-time AI services. For edge AI-empowered Internet of Things (IoT) applications, Microsoft “Azure IoT Edge”, Google “Cloud IoT”, Amazon “Web Services IoT”, and NVIDIA “EGX” provide edge AI platforms to bring real-time AI services across a wide range of applications, spanning from live video analytics [124], smart home [125], and industrial IoT [126].
III-C Lessons Learned
In the 6G era, LLMs can be deployed distributively across edge devices and servers within mobile edge networks. However, this deployment poses challenges due to the significant communication overhead and computing demands of LLMs. Therefore, it is necessary to advance MEI to efficiently support LLMs while maintaining superior performance, thereby expanding the applicability of LLMs to mobile AI applications within edge environments. As illustrated in Fig. 6, integrating LLMs into mobile edge networks leads to numerous research problems, such as model caching and delivery, training, inference, and resource management.
Figure 6: The convergence of LLMs and MEI.
IV Preliminaries II: Resource-efficient LLM Techniques
As we move forward to discuss the problems outlined in Fig. 6, we first elaborate on on-device LLM, i.e., deploying LLMs on a resource-constrained device. On-device LLM deployment, including on-device inference and training, presents the following challenges.
@@ IV Preliminaries II: Resource-efficient LLM Techniques
  • Excessive computing overhead: The computing overhead of on-device LLM deployment can be excessive. For instance, a LLAMA2-7B [80] model with 7 billion parameters requires about 1700 giga floating point operations [127] to perform a single forward pass on a sequence of 128 tokens. The Hexagon NPU, which is an advanced neural engine equipped on the Qualcomm Snapdragon X ELITE, can only provide a computing power of 45 tera operations per second. Furthermore, the backward propagation typically requires more computational resources than a forward pass [127], implying that training would be even more challenging than inference.
@@
  • Huge storage/memory demands: On the one hand, caching LLMs on edge devices consumes significant storage resources. LLMs specifically designed for on-device deployment even have billions of parameters, e.g., Google’s on-device Gemini Nano-2 has 3.25 billion parameters. On the other hand, optimizing the model with the commonly used Adam optimizer during training usually requires 12 times the memory resources needed for inference [128], which can be unacceptable for mobile devices with limited memory. These facts indicate that deploying LLMs on edge devices for both training and inference poses stringent requirements on the storage and memory resources of edge devices.
@@
  • High energy costs: The limited volume of battery in edge devices hinders the deployment of LLMs on edge devices. For instance, running an LLM quantized to INT4 with 13 billion parameters using llama.cpp (one of the most lightweight on-device LLM engines) on a Xiaomi 11 smartphone results in an energy consumption of about 56 J/token [129]. This implies that a smartphone with a battery capacity of 3000 mAh and an output voltage of 3.7 V can generate only approximately 700 tokens if the LLM is deployed on the smartphone.
@@
To mitigate the above challenges, in this section, we will review the enabling techniques tailored for resource-efficient LLM deployment, focusing on resource-efficient inference and fine-tuning. These techniques are summarized in Fig. 7, and the comparison of related works is shown in Table III. It is worth noting that the methods discussed in this section can decrease the complexity of LLM deployment on edge devices, edge servers, or in device-server collaboration. Consequently, these key techniques serve as the foundation for MEI4LLM and all subsequent sections.
@@
TABLE III: Summary of related works on resource-efficient LLM techniques.
@@
IV-A Resource-efficient Inference
On-device LLM inference eliminates privacy leakage and the need for Internet connections. However, it requires substantial computing, memory, and energy resources. In what follows, we briefly present how to mitigate these problems by introducing efficient on-device LLM inference techniques.
IV-A1 LLM compression
LLM compression enables the deployment of compressed LLMs on edge devices. The design objective is to compress LLMs without substantially compromising inference accuracy. Although compression has been extensively studied in the field of traditional DNNs, the unique architectures and properties of LLMs necessitate the redesign of existing compression strategies. The compression techniques tailored for LLMs are detailed below.
@@ IV-A1 LLM compression
In LLM quantization, uneven weight quantization can be adopted to preserve the inference accuracy of quantized LLMs since not all weights in LLMs contribute equally to the final inference results [131]. For example, the authors in [131] propose the activation-aware weight quantization for uneven weight quantization. This method quantizes most of the weights in LLMs from FP16 to INT3/INT4 while only retaining approximately 1% of the critical weights in FP16 format.
@@
IV-A2 Fast decoding for LLMs
Fast decoding methods can be used in on-device LLM deployment to save computing resources of edge devices in inference and facilitate on-device inference. As explained in Section III, LLMs typically generate tokens in an autoregressive manner, producing one token at a time based on previously generated tokens. However, as the size of LLMs grows, the volume of LLM parameters and intermediate data involved during inference also increases, leading to higher computational demands and prolonged inference latency. Additionally, as the output length increases, the LLM must repeat the decoding phase multiple times while caching all intermediate data, placing a significant burden on computing memory. To address these challenges and enable timely inference on resource-limited devices, fast decoding techniques have been developed. These techniques include leveraging lightweight LLMs for initial token generation, minimizing the volume of activated parameters/blocks, generating multiple subsequent tokens in a single forward pass, and reducing the volume of intermediate data. By applying these methods, memory footprint and inference latency can be significantly reduced. Typical fast decoding methods are elaborated below.
@@ IV-A2 Fast decoding for LLMs
Early exit: To minimize computing latency, the early exit strategy can be employed to bypass certain subsequent layers that input tokens traverse, effectively accelerating inference time. In the case of Transformers, early exit modules may be incorporated following some early blocks [137, 163], where the output from intermediate layers is converted into final outputs at exit points, provided that the desired confidence value is attained. However, the conventional early exit techniques may not be effective in LLM inference since the hidden states of the token, which output the results in an early exit layer, may be missing in the later layers [164]. For example, upon early exit, the hidden states and the KV caches of the corresponding token in the subsequent layers will be missing, hindering the generation of future token sequences [165]. To address this challenge, the hidden states of the current token can be forwarded from the exiting layer to subsequent layers for KV cache computation [138, 166] if the token outputs the result in the exiting layer.
@@
IV-B Resource-efficient Fine-tuning
@@
IV-B1 Parameter-efficient fine-tuning
@@
IV-B2 Zeroth-order optimization
Zeroth-order optimization [182] is a novel technique for model training, which estimates gradient updates only through the forward pass. This method considerably reduces the computational burden since the forward pass, equivalent to inference, demands far fewer computing resources than the backpropagation during training. Specifically, compared with prevalent first-order optimizers, such as Adam, zeroth-order optimizers do not need to store the intermediate results of the backpropagation during the training process, significantly reducing the memory consumption of LLM training. For example, the authors in [128] propose the zeroth-order optimizer, MeZO, which adopts simultaneous perturbation stochastic approximation to estimate model gradients only through the forward propagation and uses the estimated gradients to update model parameters. Compared with the fine-tuning with Adam, the model fine-tuned with MeZO demonstrates competitive performance on 7 out of 11 tasks, using only 1/12 of the running memory while only causing less than a 1% accuracy degradation. Moreover, to further enhance the efficiency of LLM fine-tuning, the zeroth-order optimization technique can be combined with PEFT methods, such as LoRA and prefix fine-tuning [128].
IV-C Lessons Learned
Resource-efficient LLM techniques enable efficient LLM training and inference on resource-limited edge devices and edge servers. Resource-efficient inference techniques focus on two main areas: model compression reduces computational demands by shrinking model sizes, whereas fast decoding improves inference speed and efficiency by optimizing data processing, reducing parameter use, and minimizing intermediate data generation. Resource-efficient fine-tuning techniques also have two primary focuses: they either shrink the number of parameters for updating or execute the forward pass only.
The overarching goal of these resource-efficient techniques is to minimize the computing workload and memory footprint, which is essential for deploying LLMs on devices with limited resources. However, there are still limitations to these approaches, as they tend to focus on single devices and do not fully exploit networked computing resources, thus placing less emphasis on communications between devices. By connecting intelligent devices, more effective model inference and training can be achieved through collaboration among edge devices and edge servers, which is the primary focus of MEI.
@@ V An Overview of MEI4LLM
As alluded to in Section IV, while there are great efforts in pushing LLMs to the network edge, the aforementioned schemes focus on implementations on a single device, limiting its applicability and efficiency. By deploying LLMs in MEI, edge devices can collaborate with edge servers for collaborative learning and inference, which largely mitigates the resource scarcity of edge devices. In line with the “NET4AI” (network for AI) vision for the 6G era [183], this section presents an overview of the MEI framework that supports the deployment of LLMs, called MEI4LLM, as shown in Fig. 9. The primary objective of MEI4LLM is to enable efficient collaboration among networked edge devices/servers with enhanced communication, computing, and storage efficiency. Clearly, this framework should also seamlessly integrate with security defense mechanisms to combat various attacks, such as data poisoning and model inversion attacks during training, evasion attacks during inference [184, 185], and unauthorized access during model and data storage, to safeguard the security of the framework. However, since the security aspect is not our main focus, we will briefly discuss it in Section IX for future research.
This section will begin by discussing the AI-native architecture specifically designed for MEI4LLM. Then, we will introduce the core components – covering parameter-sharing LLM caching and delivery, distributed LLM training (fine-tuning), and distributed LLM inference – that enable MEI4LLM to efficiently support the deployment and operations of LLMs within mobile edge networks. Further technical discussions of these core components will be elaborated in Sections VI, VII, and VIII, respectively.
@@
Figure 9: The MEI architecture for LLMs in 6G with four basic components, i.e., AI-native architecture, parameter-sharing LLM caching and delivery, distributed LLM training (fine-tuning), and distributed LLM inference.
V-A The AI-native Architecture
6G is evolving into an AI-native architecture. To effectively support pervasive LLM applications, mobile networks are expected to have the following transformative aspects: 1) Adopting the “task-oriented” design principle. Instead of maximizing throughput or minimizing latency, the design goal can be to optimize the quality of the output, say, the cross entropy of output tokens from LLMs. 2) The mobile networks must support native model partitioning across edge servers and devices to facilitate real-time LLM inference/training. 3) information-centric networking protocols can be supported to enable fast LLM delivery over the edge networks with reduced latency.
@@ V-A The AI-native Architecture
Network virtualization is of paramount importance in achieving task-oriented design. Following the design principle of software-defined networking, the MEI4LLM needs central controllers that orchestrate network-wide computing resources and data transmissions, with the decoupled control and data plane. Moreover, since the design goal and optimization principles can vary across different LLM tasks, the protocols on the control plane must be fully programmable for specific LLM services. By collecting global network knowledge, such as the accuracy of LLMs, various quantization levels, user requirements on LLM services, channel conditions, battery status of users, and computing resource availability, the controller executes the customized algorithm to coordinate model training/inference and delivery across the distributed edge computing systems, with intermediate smashed data (i.e., intermediate activations and back-propagated gradients), model parameters, or user data exchanged across edge routers and servers.
@@
Further edge networks will evolve into the “neural edge” [183], where neural network layers are distributed across edge nodes for collaborative computing. Analogous to cloud data centers with many GPUs to support large-scale LLMs, MET4LLM must feature flexible and model splitting for training and inference across distributed edge devices and servers. The optimal model splitting, placement, and data routing for large-scale models should be concertedly supported over edge networks. Moreover, both air interfaces and network designs should have native support for federated learning, split learning, and split inference for AI models, including LLMs. Since model training and inference are robust to packet errors, task-oriented wireless transmissions, say, for smashed data at the cut layer, can be carried out with appropriate error control to achieve the best efficiency-reliability trade-off.
@@
Lastly, since LLMs can be communication-intensive, information-centric networking can be implemented to ensure seamless model, feature, and data transmission across edge networks for efficient delivery of LLMs. In this respect, MEI4LLM should support LLM parameter block naming and name-based transmission protocol. By assigning names for popular LLM parameter blocks, the central controller in the MEI4LLM architecture can forward the parameter request to edge caches located nearby and support parameter block multicasting for users, thereby reducing latency and bandwidth consumption for delivering large-scale models across networks and to end users.